OpenStudy (anonymous):

Find the absolute value |1-i|

OpenStudy (anonymous):

the absolute value of \(a+bi\) is \(\sqrt{a^2+b^2}\) in your case \(a=1.b=-1\)

OpenStudy (anonymous):

I've tried -1, still saying incorrect.

OpenStudy (anonymous):

ok first off, the absolute value is always positive, so it can't be -1

OpenStudy (anonymous):

what is \(1^2\)?

OpenStudy (owen3):

$$ |1 - i | = \sqrt{ 1 + (-1)^2 } = \sqrt{ 1 + 1 }$$ can you finish it